OpenStudy (anonymous):

Okay so I'm really confused on finding percent yield I keep doing these problems and they aren't adding up can someone please check my work?

OpenStudy (anonymous):

1. Write the balanced equation for the reaction conducted in this lab, including appropriate phase symbols. Mg(s) + 2HCl(aq) --> MgCl2(aq) + H2(g) 2. Determine the partial pressure of the hydrogen gas collected in the gas collection tube. change 22.4torr to atm 22.4torr x [1atm / 760torr] = 0.0295atm. Ptotal = Pwater + Phydrogen Phydrogen = Ptotal - Pwater = 1.10atm - 0.0295 = 1.0705atm. 3. Calculate the moles of hydrogen gas collected. n = PV / RT = 1.0705atm x 0.0322L / (0.0821Latm/moleK x 297K) = 0.00141moles H2 4. If magnesium was the limiting reactant in this lab, calculate the theoretical yield of the gaseous product. Show all steps of your calculation. 0.034gMg x [1moleMg/24.3gMg] x [1moleH2 / 1moleMg] = 0.00140moles 0.00140 x 22.4L = .03136L 5. Determine the percent yield of this reaction, showing all steps of your calculation.

sam (.sam.):

Percentage yield is just \[\frac{actual~yield}{theoretical~yield} \times 100\%\]

OpenStudy (anonymous):

So it would be 0.44 % I'm just not sure if the rest of the work is correct

sam (.sam.):

Did they gave you the actual yield?

sam (.sam.):

What's the original question?

OpenStudy (anonymous):

No this is all I have plus this Mass of Magnesium Strip (grams) 0.034 Volume of gas collected (mL) 32.2 Barometric pressure (atm) 1.1 Room temperature (C ) 24 Vapor pressure of the water (torr) 22.4

sam (.sam.):

Hmm, the volume of gas collected should be the actual yield

OpenStudy (anonymous):

So then it's the moles of gas that's the actual yield?

sam (.sam.):

yeah but it seems that you're getting almost 100% yield, that's weird :\

sam (.sam.):

But if the gas are almost ideal, and the reactant is limiting reagent completely reacts and that there's no gas escaped/ any changes then you should get almost the theoretical yield

sam (.sam.):

@chmvijay

OpenStudy (anonymous):

I get 0.0998%, i converted mL to meters cubed and used R as 8.31

mathslover (mathslover):

I got 100% ( exact : 100.00003 ... %)

mathslover (mathslover):

My last equation set up was : % Yield = 2.82396 / 2.823 * 100%

sam (.sam.):

Yes that's what I'm getting as well, but something doesn't feel right, @chmvijay

OpenStudy (anonymous):

Where did you get 24.3gmL in the fourth problem